OpenStudy (anonymous):

6/y+9=5/y-9 pls help me

OpenStudy (anonymous):

y=-1/18

OpenStudy (agreene):

y=-1/18

OpenStudy (anonymous):

6/y+9=5/y-9 or, 6/y-5/y=-18 or, 1/y=-18 or, y=-1/18